Project

General

Profile

NDOS Commissioning Meeting 2011/02/21

Attendance: Jon, Zeb, Athans, Jarek, Enrique, Tricia, Alex, Minerba, Alena, Gary, Leon, Mark Mathis, Susan

Agenda:

  1. DAQ Update & Plans
  2. Channel Recovery Plans
  3. DCS Status & Plans
  4. Shifters' report
  5. AOB

Channel Recovery (Minerba):

  • 34 of 37 APDs have been replaced
  • Leon: APDs in "orange" category were marked as bad because a few pixels were noisy. Would like to try cutting ribbon cable connections for individual pixels to see if this remedies the problem.

DAQ (Leon):

  • from the DSO scans, we are now able to create a set of thresholds for every individual channel and apply them across the detector.
  • now, ~all "good" channels are running at 40-60 ADC count threshold
  • we've been running like this for most of the week
  • data are uploaded into dB, and read back out
  • Susan/Leon: we're now running with regulated voltage. Currently we're using one HV setting for all
  • Gary: what's the status of cooling? Leon: retrofit of boards is being worked on to reduce noise, should be ready for initial tests ~next week.

DCS (Athans):

  • has GUI working for DCM DCS monitoring
  • will commit code to CVS soon

Shifter report:

  • Susan:
    • DAQ group is working hard now to figure out what's going in the CR right now (can't get data flowing)
    • in general, things went pretty well this week. Documentation could be updated. In particular, the "reboot TDU" process is undocumented.

AOB:

  • doc-dB 5670: "Preliminary look of Good and Bad channels in NDOS"